# RCPU: Rotation-Constrained Error Compensation for Structured Pruning of a Large Language Model

### v1

@@ 4.3 Benchmark
Table 2: Zero-shot accuracy ($\uparrow$) on Benchmark datasets. The best score in each setting is highlighted in bold, while the second-best score is underlined.
@@
FLAP 10% 73.33 77.37 72.81 68.59 70.54 41.13 42.80 63.80
@@
FLAP 20% 70.45 74.91 67.42 67.64 66.79 39.00 43.00 61.32
@@
FLAP 30% 66.67 71.49 59.53 61.56 59.76 34.81 39.60 56.20
@@
Table 2 reports accuracy on seven language understanding benchmarks. Overall, performance degrades as the pruning ratio increases. RCPU consistently outperform WANDA-sp baseline, and the gap widens at higher pruning ratios. Moreover, RCPU achieves higher mean accuracy than FLAP across all pruning levels, indicating that geometry-preserving compensation can be more effective than bias-only correction. Looking at individual tasks, HellaSwag and WinoGrande show relatively stronger performance with RCPU. These tasks require contextual consistency and pronoun resolution, both of which are sensitive to orientation shifts in the representation space. The benefit observed here aligns with the perplexity improvements reported earlier, suggesting that rotation-constrained updates help preserve the structural properties of output representations that underlie these tasks. Similarly, for BoolQ and PIQA, which rely directly on basic language modeling ability and commonsense judgments, RCPU maintains stable performance, again consistent with trends in perplexity.
@@ 4.4 Analysis
4.4.1 Where to apply rotation?
Table 3: Perplexity under different pruning ratios and compensation targets.
@@ 4.4.1 Where to apply rotation?
In order to clarify which parts are effective to apply RCPU, we conducted ablation study that changes the module to apply the proposed compensation. Table 3 reports perplexity when rotation-based compensation is applied to different projection sub-layers. First, applying compensation to either o_proj or down_proj alone improves perplexity compared to the uncompensated baseline. This indicates that the kept subspace indeed contains recoverable signal, and aligning it to the original outputs partially restores the lost information. Second, updating o_proj is consistently more effective than updating down_proj. A plausible explanation lies in the forward order of computation in transformer blocks: attention is followed by the MLP. Misalignment at o_proj propagates directly into the subsequent MLP input, thereby amplifying its negative effect. Correcting the orientation earlier at o_proj provides the MLP with already aligned features, reducing the burden of later layers. In contrast, compensating only down_proj cannot undo the upstream misalignment originating from o_proj, and thus achieves a smaller gain. Finally, applying compensation to both o_proj and down_proj yields the largest improvement, suggesting that errors at the two sites are complementary. Moreover, the benefit becomes larger at higher pruning ratios, where the retained subspace is smaller and orientation recovery plays a more critical role.
4.4.2 Efficiency
Table 4: Size of the pruned model and time required for pruning.
@@ 5 Related Work
5.1 Model compression in LLMs
Large language models (LLMs) incur substantial computational and memory costs, motivating the development of compression techniques. Among the most widely studied approaches are quantization, which reduces parameter precision to improve efficiency Lang et al. (2024), and distillation, which transfers knowledge from a large teacher to a smaller student model Yang et al. (2024). In this work, we focus on pruning, which removes unnecessary parameters.
5.2 Unstructured pruning
Unstructured pruning accelerates inference by sparsifying weight matrices. SparseGPT Frantar & Alistarh (2023) enables one-shot pruning of LLMs via an efficient second-order update. WANDA Sun et al. (2024) introduces an activation-aware importance score that works with only a small calibration set. There also exist approaches that pursue higher accuracy, such as AlphaPruning Lu et al. (2024), which varies the pruning ratio across layers. While effective, unstructured methods do not actually reduce parameter count. Their memory and speed benefits depend on specialized sparse kernels and hardware support, and thus they are generally unsuitable for small devices.
5.3 Structured Pruning
Structured pruning removes parameters at the level of rows, columns, or blocks, directly shrinking model size and memory footprint. LLM-Pruner Ma et al. (2023) demonstrates structured pruning for LLMs but typically requires downstream fine-tuning, and Shortened-LLaMA Kim et al. (2024) prunes depth (layers) with retraining. In contrast, methods such as Wanda-sp and FLAP are applicable without additional retraining An et al. (2024): Wanda-sp extends WANDA’s activation-aware rule to column pruning, while FLAP compensates post-pruning errors via a bias term. SliceGPT Ashkboos et al. (2024) leverages a computational invariance of RMSNorm-connected transformers: by applying an orthogonal reparameterization (derived via principal component analysis), the model is reformulated into a rotated basis where entire rows and columns can be deleted while preserving functional equivalence. This is completely different from our approach, which we focus on reducing the pruning error in each layer and such update is analytically given. In contrast, our work is closer to FLAP: rather than exploiting such invariances for slicing, FLAP also tries to minimize pruning-induced output error and achieve better performance.
@@ 6 Conclusion
In this paper, we proposed RCPU, a rotation-constrained error compensation method, for structured pruning of large language models. By formulating post-pruning recovery as an Orthogonal Procrustes problem, our approach preserves the geometry of output representations while re-aligning the retained subspace to the original outputs. To complement this update, we introduced a variance-aware importance score that preferentially retains columns contributing to principal output directions, thereby enhancing the effectiveness of rotation-constrained compensation.
Through experiments on LLaMA-7B, we demonstrated that RCPU consistently reduces perplexity and improves task accuracy across pruning ratios, outperforming existing baselines such as WANDA-sp and FLAP. The improvements were particularly pronounced at higher pruning levels, indicating that geometry-preserving updates become increasingly critical as the retained subspace shrinks. Moreover, we showed the method requires no additional architectural changes and only modest computational overhead, making it practically applicable to large-scale deployments.
Overall, our findings highlight the importance of incorporating geometric constraints into error compensation for pruning. We believe that the proposed framework opens up new directions for designing pruning-aware model updates that are both statistically stable and computationally efficient, and it can serve as a foundation for further advances in scalable and reliable model compression.

### v2

@@ 4.3 Benchmark
Table 2: Zero-shot accuracy ($\uparrow$) on Benchmark datasets when $N_{\mathrm{calib}}=128$ on Llama-7B. The best score in each setting is highlighted in bold, while the second-best score is underlined.
@@
Prop.Score+LS ($\lambda_{\mathrm{best}}$) 10% 72.75 75.19 71.03 68.19 64.65 39.93 39.60 61.62
@@
Prop.Score+LS ($\lambda_{\mathrm{best}}$) 20% 67.83 71.71 64.40 64.56 58.71 34.81 35.80 56.83
@@
Prop.Score+LS ($\lambda_{\mathrm{best}}$) 30% 62.54 67.08 55.55 60.77 50.13 31.06 35.00 51.73
@@
Table 3: Zero-shot accuracy ($\uparrow$) on Benchmark datasets when $N_{\mathrm{calib}}=512$ on Llama-2-13B.
Method P.R. BoolQ PIQA Hella Wino ARC-e ARC-c OBQA Mean
Llama2-13B (Orig.) 0% 80.55 79.05 79.37 72.14 77.44 49.06 45.20 68.97
FLAP 10% 74.22 78.56 76.12 71.11 74.07 44.54 45.20 66.26
SliceGPT 10% 62.84 77.09 71.80 71.59 76.35 49.40 45.20 64.90
WANDA-sp 10% 79.14 78.02 77.99 70.64 75.76 48.12 44.80 67.78
Prop.Score+LS ($\lambda_{\mathrm{best}}$) 10% 78.78 77.69 77.56 72.30 73.99 47.53 43.40 67.32
RCPU (Rot.) 10% 79.82 78.13 78.09 72.45 75.00 47.78 44.40 67.95
RCPU(Rot.+Scale) 10% 79.79 78.29 78.14 72.69 75.04 47.78 44.60 68.05
FLAP 20% 67.00 74.97 70.41 68.19 67.09 40.78 43.20 61.66
SliceGPT 20% 52.20 71.76 63.17 67.32 70.45 43.77 41.80 58.64
WANDA-sp 20% 72.78 76.61 73.32 69.46 72.39 44.80 41.80 64.45
Prop.Score+LS ($\lambda_{\mathrm{best}}$) 20% 73.12 76.28 73.02 69.93 70.50 43.86 41.20 63.99
RCPU (Rot.) 20% 73.76 76.44 73.91 70.09 71.25 44.20 42.20 64.55
RCPU (Rot.+Scale) 20% 73.30 76.33 73.95 69.46 71.21 43.34 41.80 64.20
FLAP 30% 65.78 72.14 64.57 64.25 62.71 38.91 40.20 58.37
SliceGPT 30% 38.35 66.10 52.64 66.38 56.78 35.15 40.00 50.77
WANDA-sp 30% 61.99 62.68 36.09 51.07 41.54 25.60 28.80 43.97
Prop.Score+LS ($\lambda_{\mathrm{best}}$) 30% 66.61 72.96 64.62 65.04 66.62 36.77 40.40 59.00
RCPU (Rot.) 30% 64.37 73.72 66.22 64.88 67.05 38.31 42.80 59.62
RCPU (Rot.+Scale) 30% 65.17 73.67 66.69 64.88 67.30 38.23 42.60 59.79
Table 2 and Table 3 reports accuracy on seven language understanding benchmarks on Llama-7B and Llama-2-13B. Overall, performance degrades as the pruning ratio increases. RCPU achieves higher mean accuracy than FLAP across all pruning levels, indicating that geometry-preserving compensation can be more effective than bias-only correction. Comparing Table 2 with Table 6, the scaled variant performed better and ranked as the best baseline more often in the 512-sample setting than in the 128-sample setting. Intuitively, the additional samples stabilize the norm statistics of the unpruned versus pruned outputs, allowing the global scale $s^{\star}$ to more effectively restore the original magnitude. Regarding Table 3, SliceGPT transforms the entire model into an equivalent structure using an orthogonal matrices, and then performs row or column deletion in a single global step. Due to this property of applying a global transformation followed by global pruning, we expect that, at high pruning ratios, a mismatch arises between the input distributions assumed by each layer and the actual input distributions after pruning. This mismatch is also expected to accumulate across layers and thus we believe accuracy tends to degrade at high pruning ratios. In contrast, RCPU optimizes the pruning-induced error layer by layer. Thus, even at high pruning ratios, each layer is more likely to maintain representations close to the inputs it assumes, which we believe leads to better benchmark performance. Looking at individual tasks, HellaSwag and WinoGrande show relatively stronger performance with RCPU. These tasks require contextual consistency and pronoun resolution, both of which are sensitive to orientation shifts in the representation space. The benefit observed here aligns with the perplexity improvements reported earlier, suggesting that rotation-constrained updates help preserve the structural properties of output representations that underlie these tasks.
@@ 4.4 Analysis
Where to apply rotation?
Table 4: PPL under different pruning ratios and compensation targets (Llama-7B, $N_{\mathrm{calib}}=128$).
@@
In order to clarify which parts are effective to apply RCPU, we conducted ablation study that changes the module to apply the proposed compensation. Table 4 reports PPL when rotation-based compensation is applied to different projection sub-layers. First, applying compensation to either o_proj or down_proj alone improves PPL compared to the uncompensated baseline. This indicates that the kept subspace indeed contains recoverable signal, and aligning it to the original outputs partially restores the lost information. Second, updating o_proj is consistently more effective than updating down_proj. A plausible explanation lies in the forward order of computation in transformer blocks: attention is followed by the MLP. Misalignment at o_proj propagates directly into the subsequent MLP input, thereby amplifying its negative effect. Correcting the orientation earlier at o_proj provides the MLP with already aligned features, reducing the burden of later layers. In contrast, compensating only down_proj cannot undo the upstream misalignment originating from o_proj, and thus achieves a smaller gain. Finally, applying compensation to both o_proj and down_proj yields the largest improvement, suggesting that errors at the two sites are complementary. Moreover, the benefit becomes larger at higher pruning ratios, where the retained subspace is smaller and orientation recovery plays a more critical role.
Efficiency
Table 5: Size of the pruned model and time required for pruning (Llama-7B).
@@ 5 Related Work
Model compression in LLMs
Large language models (LLMs) incur substantial computational and memory costs, motivating the development of compression techniques. Among the most widely studied approaches are quantization, which lowers precision to improve efficiency Lang et al. (2024), and distillation, which transfers knowledge from a large teacher to a smaller student model Yang et al. (2024). In this work, we focus on pruning, which removes unnecessary parameters.
Unstructured pruning
Unstructured pruning accelerates inference by sparsifying weight matrices. SparseGPT Frantar and Alistarh (2023) enables one-shot pruning of LLMs via an efficient second-order update. WANDA Sun et al. (2024) introduces an activation-aware importance score that works with only a small calibration set. There also exist approaches that pursue higher accuracy, such as AlphaPruning Lu et al. (2024), which varies the pruning ratio across layers. While effective, unstructured methods do not actually reduce parameter count. Their memory and speed benefits depend on specialized hardware supports, and thus they are generally unsuitable for small devices.
Structured Pruning
Structured pruning removes parameters at the level of rows, columns, or blocks, directly shrinking model size and memory footprint. LLM-Pruner Ma et al. (2023) demonstrates structured pruning for LLMs but typically requires downstream fine-tuning, and Shortened-LLaMA Kim et al. (2024) prunes depth (layers) with retraining. In contrast, methods such as Wanda-sp and FLAP are applicable without additional retraining An et al. (2024): Wanda-sp extends WANDA’s activation-aware rule to column pruning, while FLAP compensates post-pruning errors via a bias term. SliceGPT Ashkboos et al. (2024) leverages a computational invariance of RMSNorm-connected transformers: by applying an orthogonal reparameterization (derived via principal component analysis), the model is reformulated into a rotated basis where entire rows and columns can be deleted while preserving functional equivalence. Rotation has also been used to improve prunability in other forms: RotPruner Chen and Wang (2025) learns layer-wise orthogonal transforms to obtain pruning-friendly parameterizations, and DenoiseRotator Gu et al. (2025) trains rotations that concentrate importance scores before pruning. These methods use rotation as an additional parameterization to facilitate pruning before or during parameter removal. Unlike these rotation-learning or slicing-based approaches, RCPU focuses on reducing the pruning error after structured column removal through an analytically derived orthogonal compensation. Our work is closer to FLAP, trying to minimize pruning-induced output error and achieve better performance.
@@ 6 Conclusion
In this paper, we proposed RCPU, a rotation-constrained error compensation method, for structured pruning of large language models. By formulating post-pruning recovery as an Orthogonal Procrustes problem, our approach aimed to preserve the geometry of output representations while re-aligning the retained subspace to the original outputs. To complement this update, we adopted a variance-aware importance score that preferentially retains columns contributing to principal output directions, thereby enhancing the effectiveness of rotation-constrained compensation. Through experiments on Llama-7B and Llama-2-13B, we demonstrated that RCPU improves perplexity and task accuracy across pruning ratios, outperforming existing baselines. The improvements were particularly pronounced at higher pruning levels, indicating that geometry-preserving updates become increasingly critical as the retained subspace shrinks. Moreover, we showed the method requires no additional architectural changes and only modest computational overhead, making it practically applicable to large-scale deployments.
Overall, our findings suggest that incorporating geometric constraints into error compensation for pruning can be a promising direction. We hope that this framework stimulates further investigation into pruning-aware model updates that balance statistical stability and computational efficiency.
Ethics and Reproducibility Statement
Disclosure of AI assistance.
We used a large language model to edit and polish the manuscript text. All research ideas, methods, and experiments were conducted solely by the authors.
Data usage and privacy.
All calibration and evaluation datasets used in this work are publicly available and contain no sensitive personal information. We used the data under their respective licenses, made no attempts at re-identification, and did not store or share model inputs and outputs beyond the scope of calibration and evaluation.
Environmental impact.
This study adds only a small incremental compute footprint: calibration consists of forward passes plus one small SVD per targeted sub-layer. We did not perform gradient-based fine-tuning in our experiments. At deployment time, structured pruning reduces parameter count and effective FLOPs, which can lower inference cost under comparable hardware and batching conditions.
Fairness and safety.
Structured pruning can alter performance unevenly across tasks, domains, or languages. We therefore evaluate on diverse benchmarks and report per-task metrics to surface potential regressions. No safety-critical deployment is claimed.
Reproducibility.
All experimental settings and tools are described in the main text, including the models and datasets and the calibration setup.